Chris invested $4,100 into two money market accounts. One account has an interest rate of 3% and the other has a rate of 7%. Total interest earned at the end of the year was $191. How much was invested in each account?
Answer by stanbon(75887)

Chris invested $4,100 into two money market accounts. One account has an interest rate of 3% and the other has a rate of 7%. Total interest earned at the end of the year was $191. How much was invested in each account?
----
Equation:
int + int = 191
----
0.03*x + 0.07(4100-x) = 191
-------
3x + 7*4100 - 7x = 19100
-----------
-4x = -9600
x = $2400
----
4100-2400 = $1700
-------------
Cheers,
Stan H.
==============